OpenStudy (sammietaygreen):

which polynomial function best models the set data below. x 0 2 4 6 8 10 12 14 16 18 20 f(x) 2 5 5 4 2 0 -2 -2 0 5 14

OpenStudy (sammietaygreen):

I think I have the correct answer choice, just not positive.

OpenStudy (maddy1251):

Do you have some functions?

OpenStudy (maddy1251):

When f(x) is f(2) it should equal 0, that is when you plug in '2' into your equaltion. e.g f(x)=x-2, so f(2)=2-2 = 0 . Only because it states x=0 when f(x)=f(2).

imqwerty (imqwerty):

what are the options

OpenStudy (sammietaygreen):

@Maddy1251 that's what I did

OpenStudy (maddy1251):

@sammietaygreen all you gotta do is plug in the f(x) options into the equations.

OpenStudy (sammietaygreen):

Is my answer choice correct though?

OpenStudy (superdavesuper):

yes u got it @sammietaygreen

OpenStudy (sammietaygreen):

thank you!

OpenStudy (superdavesuper):

welcome - u did the work urself anyway :)